# IndicMMLU-Pro: Benchmarking the Indic Large Language Models

### v1

@@ 5 Key Findings and Implications of IndicMMLU-Pro
IndicMMLU-Pro represents a significant step forward in the development and evaluation of AI models for the Indic languages. Through this comprehensive benchmark, we have shed light on the current capabilities and limitations of state-of-the-art multilingual models in handling the linguistic diversity and complexity of the Indian subcontinent. This section details implications based on some of our key findings.
@@ 5.1.1 GPT-4o Dominance
GPT-4o consistently outperformed all other models across all languages, with accuracy scores ranging from 38.46% to 44.80%. This suggests that large, advanced language models have significant potential for handling the Indic languages.
@@ 5.1.2 Performance Tiers
While GPT-4o and GPT-4o mini showed superior performance, specialized models like XLM-RoBERTa and Navarasa consistently outperformed other models in the foundation tier. This highlights the importance of both scale and specialized training for the Indic languages.
@@ 5.4.3 Evaluation Metrics
These findings and implications highlight the significant advancements made by IndicMMLU-Pro in evaluating and understanding the performance of language models across the Indic languages, while also pointing to crucial areas for future research and development in multilingual NLP.
@@ 5.5 Future Directions
  • Data Collection: There is a pressing need for more high-quality, diverse datasets across all the Indic languages, particularly for low-resource languages. This will in turn enable more robust model training and evaluation.
  • Model Development: Future research should focus on developing models that can better handle the unique linguistic features of the Indic languages, including their complex morphology and script diversity. This may involve innovative architecture designs or novel pre-training techniques.
  • Cross-lingual Transfer: Exploring techniques to improve knowledge transfer between related the Indic languages will help boost performance, especially for low-resource languages. This could involve leveraging linguistic similarities or developing more effective multilingual training strategies.
@@
  • Evaluation Metrics: Further refinement of evaluation metrics to account for the linguistic and cultural nuances of the Indic languages is necessary for more accurate performance assessment. Existing metrics often overlook the linguistic and cultural nuances inherent to the Indic languages.
IndicMMLU-Pro sets a new standard for evaluating AI models in the context of the Indic languages. By providing a comprehensive, multi-task, and multilingual benchmark, it enables researchers and developers to assess and improve the capabilities of their models across a wide range of indic languages and domains. As the field of NLP continues to advance, we hope that IndicMMLU-Pro will catalyze increased research and development in the technologies based on the Indic languages. This benchmark not only highlights the current state of the art but also points the way forward for creating more linguistically diverse, culturally sensitive, and capable AI systems that can serve the vast and diverse population of the Indian subcontinent.
To foster a more comprehensive understanding of the Indic language models, it is essential to intentionally integrate multi-disciplinary perspectives into engineering and science workstreams Bhaduri et al. (2024b). Collaborations between engineers, scientists, and experts from social sciences can provide nuanced insights into the complex interplay between language, culture, and technology Mackenzie et al. (2024). By embracing interdisciplinary approaches and refining evaluation metrics, we can work towards a more inclusive and accurate assessment of the Indic language models, ultimately driving progress in natural language processing for diverse languages and communities.
In conclusion, while significant challenges remain in the field of the Indic language NLP, IndicMMLU-Pro provides a robust framework for measuring progress and guiding future research efforts. As we continue to push the boundaries of what’s possible in multilingual AI, benchmarks like IndicMMLU-Pro will play a crucial role in ensuring that technological advancements benefit all language communities, fostering greater inclusivity and accessibility in the digital age.
@@ 6 Acknowledgements
We extend our deepest gratitude to the experts who volunteered their time and effort to evaluate our dataset. Unfortunately, the Indic languages are often neglected in the development and evaluation of Large Language Models, making initiatives like ours reliant on the generosity and expertise of individuals like those listed below. The list of experts and the language they reviewed are:
@@
Expert feedback has significantly enhanced the quality of our dataset, and we are grateful for their dedication to promoting the Indic languages in the global LLM community.

### v2

@@ 5 Key Findings and Implications of IndicMMLU-Pro
IndicMMLU-Pro represents a significant step forward in the development and evaluation of AI models for Indic languages. Through this comprehensive benchmark, we have shed light on the current capabilities and limitations of state-of-the-art multilingual models in handling the linguistic diversity and complexity of the Indian subcontinent. This section details implications based on some of our key findings.
@@ 5.1.1 GPT-4o Dominance
GPT-4o consistently outperformed all other models across all languages, with accuracy scores ranging from 38.46% to 44.80%. This suggests that large, advanced language models have significant potential for handling Indic languages.
@@ 5.1.2 Performance Tiers
While GPT-4o and GPT-4o mini showed superior performance, specialized models like XLM-RoBERTa and Navarasa consistently outperformed other models in the foundation tier. This highlights the importance of both scale and specialized training for Indic languages.
@@ 5.4.3 Evaluation Metrics
These findings and implications highlight the significant advancements made by IndicMMLU-Pro in evaluating and understanding the performance of language models across Indic languages, while also pointing to crucial areas for future research and development in multilingual NLP.
@@ 5.5 Future Directions
  • Data Collection: There is a pressing need for more high-quality, diverse datasets across all Indic languages, particularly for low-resource languages. This will in turn enable more robust model training and evaluation.
  • Model Development: Future research should focus on developing models that can better handle the unique linguistic features of Indic languages, including their complex morphology and script diversity. This may involve innovative architecture designs or novel pre-training techniques.
  • Cross-lingual Transfer: Exploring techniques to improve knowledge transfer between related Indic languages will help boost performance, especially for low-resource languages. This could involve leveraging linguistic similarities or developing more effective multilingual training strategies.
@@
  • Evaluation Metrics: Further refinement of evaluation metrics to account for the linguistic and cultural nuances of Indic languages is necessary for more accurate performance assessment. Existing metrics often overlook the linguistic and cultural nuances inherent to Indic languages.
IndicMMLU-Pro sets a new standard for evaluating AI models in the context of Indic languages. By providing a comprehensive, multi-task, and multilingual benchmark, it enables researchers and developers to assess and improve the capabilities of their models across a wide range of Indic languages and domains. As the field of NLP continues to advance, we hope that IndicMMLU-Pro will catalyze increased research and development of technologies based on Indic languages. This benchmark not only highlights the current state of the art but also points the way forward for creating more linguistically diverse, culturally sensitive, and capable AI systems that can serve the vast and diverse population of the Indian subcontinent.
To foster a more comprehensive understanding of Indic language models, it is essential to intentionally integrate multi-disciplinary perspectives into engineering and science workstreams Bhaduri et al. (2024b). Collaborations between engineers, scientists, and experts from social sciences can provide nuanced insights into the complex interplay between language, culture, and technology Mackenzie et al. (2024). By embracing interdisciplinary approaches and refining evaluation metrics, we can work towards a more inclusive and accurate assessment of Indic language models, ultimately driving progress in natural language processing for diverse languages and communities.
In conclusion, while significant challenges remain in the field of Indic language NLP, IndicMMLU-Pro provides a robust framework for measuring progress and guiding future research efforts. As we continue to push the boundaries of what’s possible in multilingual AI, benchmarks like IndicMMLU-Pro will play a crucial role in ensuring that technological advancements benefit all language communities, fostering greater inclusivity and accessibility in the digital age.
@@ 6 Acknowledgements
We extend our deepest gratitude to the experts who volunteered their time and effort to evaluate our dataset. Unfortunately, Indic languages are often neglected in the development and evaluation of Large Language Models, making initiatives like ours reliant on the generosity and expertise of individuals like those listed below. The list of experts and the language they reviewed are:
@@
Expert feedback has significantly enhanced the quality of our dataset, and we are grateful for their dedication to promoting Indic languages in the global LLM community.